A mini-fridge needs to maintain a temperature of 277 K. The air outside the mini-fridge has a temperature of 303 K. Every fifteen minutes, 2.88
Answer #1

1a. Power output of compressor=[Heat transferred out of mini fridge-heat transferred into surrounding]/time

=[3.6*105-2.88*105]/[15*60]=80J/sec

2a. COP= [Heat transferred out of refrigerator]/[Heat transferred out of mini fridge-heat transferred into surrounding]

=2.88*105/[3.6*105-2.88*105]

=4

3a. COPrefrigerator=TH/TH-Tc=303/[303-277]=11.65

4a. Heat transferred out= COP*Power of compressor*time in sec

=11.65*80*15*60=8.39*105 J

1b. COP=TH/[TH-TC]=302/[302-291]=27.45

2b. Heat transferred into the house= Work done*(COP+1)=1300*(27.45+1)=36.990*103 J

3b. Heat transferred into the house= Work done*(COP+1)=1300*[4+1]=6.5*103 J

1C. COP=TH/[TH-TC]=302/[302-276]=11.62

2c. Heat transferred into the house= Work done*(COP+1)=1300*(11.62+1)=16.4*103 J

3c. Heat transferred into the house= Work done*(COP+1)=1300*[4+1]=6.5*103 J